You need to prepare a paraformaldehyde (PFA) solution in order to fix cells for staining. You begin with a 32% paraformaldehyde solution. The protocol asks for a 4% paraformaldehyde fixation. If you want to make 15mL of 4% paraformaldehyde how much 32% paraformaldehyde needs to be added to what volume of PBS
Answer #1

C1V1= C2V2

C1= concentration of the stock solution

V1 - the volume of stock solution that is used to make the working solution

C2 - concentration of the working solution.

V2 - Final volume

For the given problem

C1 = 32%

V1= need to find out

C2 = 4%

V2 = 15ml

32 % * V1 = 4 % * 15

V1 = 4*15 / 32 = 1.875 ml.

The volume of PBS = 15 - 1.875 = 13.125ml.

So we need to add 1.875ml form 32% paraformaldehyde solution and 13.125 ml PBS to make 15 ml of 4% paraformaldehyde solution.

Please write to me if you need more help.